OpenStudy (anonymous):

Nakim simplified 3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x and got -10x times the square root of 2x for an answer. Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression. (You can use "sqrt()" to show a square root. For example, 7 times the square root of a can be written as 7sqrt(a). )

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

^

terenzreignz (terenzreignz):

Exemplary patience :)

OpenStudy (ivettef365):

3\(\sqrt{2x}\) + x \(\sqrt{8x}\) - 5 \(\sqrt{18x}\) = -10x\(\sqrt{2x}\)

OpenStudy (ivettef365):

that is equal to 3 \(\sqrt{2x}\) + 2x \(\sqrt{2x}\) -15 \(\sqrt{2x}\)

OpenStudy (ivettef365):

what was done wrong is that Nakim added the 3+2-15 = -10 and you can not add the 2x because of the variable, can only be added the 3 -15

OpenStudy (ivettef365):

so the answer on this would be -12\(\sqrt{2x}\) +2x \(\sqrt{2x}\)

OpenStudy (anonymous):

im sorry but can you show me how you got the answer

OpenStudy (ivettef365):

I showed you, where is the part you don't understand so I can explain

OpenStudy (anonymous):

how sid you get the answer

OpenStudy (ivettef365):

?

OpenStudy (anonymous):

im not good with these so can you possibly show me step by step how you got the answer

OpenStudy (ivettef365):

3 \(\sqrt{2x}\) + x \(\sqrt{4.2x}\) -5\(\sqrt{9.2x}\)

OpenStudy (ivettef365):

are you good up to there

OpenStudy (anonymous):

I got it Nakim started out correct but on the last step he combined all three terms when only the first and last terms can be combined. The middle term, 2x*SQRT(2x), cannot be combined with the other two. 2. 3SQRT(2x) + xSQRT(8x) - 5SQRT(18x) = 3SQRT(2x) + xSQRT(4)SQRT(2x) - 5SQRT(9)SQRT(2x) = 3SQRT(2x) + x*2SQRT(2x) - 5*3SQRT(2x) = 3SQRT(2x) + 2xSQRT(2x) - 15SQRT(2x) = -12SQRT(2x) + 2xSQRT(2x) or [SQRT(2x)]*[-12+2x]